OpenStudy (anonymous):

A superhero recently asked his nemesis how many cats she has. She answered with a riddle: "Two-thirds of my cats plus three." How many cats does she have?

OpenStudy (anonymous):

i got 11/3 idk if its right though

OpenStudy (anonymous):

This question is driving me nuts haha

OpenStudy (radar):

Where did all that extra info come from @clara.university I read only "Two-thirds of cats plus three."

OpenStudy (anonymous):

yah same here

OpenStudy (radar):

Come on GrooveThumper, how is the problem really worded?

OpenStudy (anonymous):

That's exactly how it was worded, verbatim.

OpenStudy (radar):

Where did clara.university get the info that was posted in her solution.

OpenStudy (anonymous):

maybe she made it up

OpenStudy (anonymous):

No idea.

OpenStudy (texaschic101):

3/4C + 3 = C 3 = C - 3/4C 3 = 4/4C - 3/4C 3 = 1/4C 3/(1/4) = C 3 * 4/1 = C 12 = C I think there is 12 cats

OpenStudy (anonymous):

makes sense to me ^

OpenStudy (radar):

Would you substitute the 2/3 posted in the problem for where you have 3/4?

OpenStudy (texaschic101):

wait...I wrote it wrong 2/3C + 3 = C 3 = C - 2/3C 3 = 3/3C - 2/3C 3 = 1/3C 3/(1/3) = C 3 * 3/1 = C 9 = C I found another problem like it using different numbers...sorry, I wrote the wrong problem. This is the right one.

OpenStudy (radar):

Thanks texaschic101, 9 cats it is.

OpenStudy (anonymous):

Thank you so much, Texas! I accidentally hit the "best response" on another one - otherwise I would have given it to you!

OpenStudy (texaschic101):

no worries....I am not worried about medals

OpenStudy (anonymous):

Still deserve it though! :) Thank you, again.

OpenStudy (texaschic101):

sure thing :)

OpenStudy (radar):

I gave her one, it was worth it to get rid of all that anxiety. lol

OpenStudy (anonymous):

Hahaha

OpenStudy (radar):

Plus it was a good exercise in algebra, step by step. I have one cat and he is enough!.

OpenStudy (radar):

Welcome to Open Study @GrooveThumper I believe that was your first question.

OpenStudy (anonymous):

yo @radar i got a hard one posted think you can help?

OpenStudy (anonymous):

no ones given me a response :(

OpenStudy (radar):

I go look and see.

OpenStudy (anonymous):

thanks bud :D